# Benchmarking Quantum Computer Simulation Software Packages

### v1

@@ Introduction
The rapidly growing interest in quantum computers from both the academic and commercial side fuels the need for software packages that can simulate their operation and aid both hardware and algorithm development. These software environments generally model the quantum state evolution of an isolated set of qubits acted on by a set of well-defined quantum gate operations with an intrinsic error rate set only by the classical computer’s numerical precision. Some packages further support an open quantum systems simulation, which can also model external perturbations, thereby allowing for the simulation of actual, noisy quantum hardware at the expense of additional computational overhead. In all cases, the available computational hardware, the underlying mathematical formalism [1] and its particular software implementation place stringent limits on the system sizes that can be explored. With the exception of a few special cases, even High Performance Computing (HPC) platforms often cannot reach beyond a system size of about 50 simulated qubits before they run into resource limitations [2, 3, 4, 5]. Yet, even at moderate scales of a few tens of qubits, efficient simulation with reasonable execution times is a formidable challenge due to the vast amounts of data that have to be processed to faithfully store and manipulate a generic quantum state. With an amount scaling exponentially with $2^{\textrm{\#qubits}}$, memory and the transfer of data between processing and storage are at the core of the slow-down in computation speed and this is where HPC architecture and a tailored software implementation can make a difference.
@@
qiskit[7], cirq[8], qsimcirq[9], pennylane[10], pennylane lightning[10], qibo[11], qibojit[12], Intel QS[13], projectq[14], qrack[15], qpanda[16], qcgpu[17], quest[18], qulacs[19], qpp[20], SV-Sim[21], Yao[22], HiQ[23], HybridQ[24], Braket, myQLM, QuTiP[25], PyQuil[26], pytket[27], Microsoft QDK Simulator, Blueqat (qaqarot), Quantastica Toaster, HyQuas[28], MPIQulacs[29], JUQCS[30], Quimb[31], NVidia cuQuantum[32], Spinoza[33], QuantumFlow, QPlayer[34], Torchquantum[34], pyqtorch, QCompute,QX Simulator, Basiq[35], MIMIQ (QPerfect), Qristal, QCLAB++[36]
@@
qiskit, cirq, pennylane, qibo, quantum-gates [37], QuantumSim, NWQSim [21, 38], QuaC, QuTiP (PIQS), OpenQuantumTools.jl [39]
@@
qiskit, cirq, qrack, Stim [51], QuantumClifford.jl, PyZX [52], MIMIQ (QPerfect)
@@
First, there are statevector-based simulators that simulate the evolution of pure quantum states. Here, a statevector of $N$ qubits is represented by a $2^{N}$-sized vector of complex valued entries which are modified under the action of a formally $2^{N}\times 2^{N}$-sized unitary matrix operator. A more complex (and resource intensive) simulation is offered by the density matrix formalism, which further allows to simulate mixed states, which usually arise under the influence of noise and in the case of open quantum systems. Tensor networks offer an alternate, yet approximate, representation of the quantum state. In the example of Matrix Product States (MPS) or Projected Entangled Pair States (PEPS) [67, 68] the wavefunctions are expressed in a compressed way using a set of local tensors to be contracted. The unitaries are represented as Matrix Product Operators (MPO) or Projected Entangled Pair Operators (PEPO), which are smaller in size than the above unitaries. However, with increasing entanglement in the system, the size of the tensors, referred to as the so-called bond dimension, is forced to grow to maintain an accurate approximation. Clifford algebra based simulators have been developed [69] that allow the access of qubit numbers on the order of a million and more. They are mainly used to investigate quantum error correction codes [51] and more recently random unitary circuits. Yet, these have been shown to not be universal for quantum computation [69]. Further, there are simulators that are tailored to a specific hardware architecture or application domain, aiming to emulate the corresponding platform exactly or being optimized for a specific use-case.
Table 2: Curated list of quantum simulation packages whose performance is benchmarked
Package Language Version Singlethread Multithread GPU MPI Multi-GPU
SP
DP
SP
DP
SP
DP
SP
DP
$\text{Braket}^{[k]}$
Python
1.38.1
$\checkmark$
$\checkmark$
Cirq[8]
Python
1.1.0
$\checkmark$
$\checkmark$
$\text{cuQuantum }^{[j]}$[70]
Python
22.11.0
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
HiQ[23]
Python
0.0.1
$\checkmark$
$\checkmark$
$\checkmark^{[b]}$
HybridQ[24]
Python
0.8.2
$\checkmark$
$\checkmark$
$\checkmark$
Intel-QS(cpp)[13]
C++
2.1.0
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark^{[d]}$
myQLM (py)
Python
1.7.3
$\checkmark$
@@
C++
0.0.5
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark^{[l]}$
$\checkmark^{[l]}$
Pennylane(py)[10]
Python
0.28.0
$\checkmark$
$\checkmark$
Pennylane(cpp)[10]
C++
0.28.2
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark^{[c]}$
$\checkmark^{[b]}$
Projectq[14]
Python
0.8.0
$\checkmark$
$\checkmark$
Qcgpu[17]
Python
0.1.1
$\checkmark$
Qibo[11]
Python
0.1.11
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Qibojit[12]
Python
0.0.7
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Qiskit[7]
Python
0.41.0
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark^{[a]}$
$\checkmark$
$\checkmark$
QPanda[16]
Python
3.7.16
$\checkmark$
$\checkmark$
$\checkmark^{[h]}$
Qrack[15]
C++
8.2.2
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Qsimcirq[9]
Python
0.15.0
$\checkmark$
$\checkmark$
$\checkmark^{[b]}$
$\checkmark$
Quest[18]
C
3.5.0
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark^{[d]}$
Qulacs[19]
Python
0.5.7
$\checkmark$
$\checkmark$
$\checkmark^{[e]}$
$\checkmark^{[f]}$
Q++[20]
C++
4.0.1
$\checkmark$
$\checkmark$
SV-Sim[21]
Python
1.0
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark^{[d]}$
$\checkmark^{[g]}$
Yao[22]
Julia
0.8.6
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark^{[d]}$
$\checkmark^{[d]}$
@@
In this manuscript, we restrict our analysis to statevector simulators. Furthermore, our selection is based on additional criteria, such as, whether a package offers to exploit the computational power provided by HPC platforms, is actively maintained and developed (gauged in terms of release cycles) and provides access to relevant documentation (feature snippets, examples and possibly use-case tutorials). Table 2 presents our down-selected list of packages benchmarked in this work and their supported HPC capabilities [^1] While we list the MPI capability, we do not present MPI benchmarks in this work..
@@
Figure 1: Workflow of the toolchain which takes four inputs: (i) Quantum Task, (ii) Simulation package, (iii) Compute capability of the HPC, (iv) Float precision. The task is input as the QASM instruction set which is processed by the translator to further generate the package specific instruction set. Post translation, other auxiliary file generation is triggered that creates the necessary job scripts, linker files that scope the translated run files into the environment of the container. In addition, a script to trigger the job submission onto the cluster is also generated. Finally, the performance characteristics either corresponding to the time or memory are collected for further analysis. ($*{}^{*}$) The translated files generated from OpenQASM are executed using the package installed on the container, while the pre- and post-processing steps are executed on the HPC node independent of the container.
@@
One of the first experiments that established claims of quantum advantage was the Random Quantum Circuits (RQC) sampling experiment implemented on the Sycamore chip [81]. In our benchmarks, we have used the QASM files used in the original experiment (specifically, the set marked as the EFGH pattern).
@@
QFT maps the state |$\ket{\psi}=\sum\limits_{i=0}^{N-1}\alpha_{i}\ket{i}$ where $N=2^{n}$ bitstrings (basis states) to |$\ket{\phi}=\sum\limits_{i=0}^{N-1}\beta_{i}\ket{i}$ where $\beta_{k}=\frac{1}{\sqrt{N}}\sum\limits_{n=0}^{N-1}\alpha_{n}\omega_{N}^{nk}$ with $k=0,1,2,3,...,N-1$ and $\omega_{N}=e^{\frac{2\pi i}{N}}$. The corresponding unitary is given by
$QFT_{n}=\begin{bmatrix}1&1&1&1&\dots&1\\ 1&\omega_{N}&\omega_{N}^{2}&\omega_{N}^{3}&\dots&\omega_{N}^{N-1}\\ 1&\omega_{N}^{2}&\omega_{N}^{4}&\omega_{N}^{6}&\dots&\omega_{N}^{2(N-1)}\\ 1&\omega_{N}^{3}&\omega_{N}^{6}&\omega_{N}^{9}&\dots&\omega_{N}^{3(N-1)}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&\omega_{N}^{N-1}&\omega_{N}^{2(N-1)}&\omega_{N}^{3(N-1)}&\dots&\omega_{N}^{(N-1)(N-1)}.\\ \end{bmatrix}$
@@
$QFT(\ket{\varphi})$ $=\frac{1}{\sqrt{N}}\Big{\{}\bigg{(}\ket{0}+e^{2\pi i[0.\varphi_{n-1}]}\ket{1}\bigg{)}$
@@
where $[0.\varphi_{0}\varphi_{1}...\varphi_{l}]=\sum\limits_{m=0}^{l}\varphi_{m}2^{-m}$. The $\sigma_{z}$ expectation value of Eq. (9) at any site $i$ will be equal to zero, which we use to validate the computations and order the benchmarked packages based on their precision via their proximity to zero, as detailed in a later section.
@@ Results
In this part of our study, we configure the toolchain to limit the available compute capability to a single computational thread. The aim is to investigate the effect of numerical precision (single vs. double) on the both wall-clock time scaling as well as the maximum system size that can be simulated. While this difference of 32 bit vs. 64 bit representation will not allow for a much larger number of qubits (a factor two reduction in total memory), less data will have to be processed at every step, allowing for speed-ups depending on the package internals.
@@
We start by noting that in the top row of Figs. 5 and 6, we observe that some packages exhibit almost straight lines for the entire range of system sizes, while other packages have a rather small slope at lower numbers of qubits and only cross over to the steeper slope for qubit numbers of about 20. This leads to the paradoxical situation that the fastest package in Fig. 6 at the largest qubit numbers (qiskit) is among the slowest packages for eight qubits  [^2] A similar observation holds for hybridq in the single precision case, although it is never the fastest package, even though it gets quite close..
Among the packages that support single precision, shown in Fig. 5, we can clearly identify that the package qsimcirq performs best in the large-$N$ limit. For double precision resolution, shown in Fig. 6, the best performing package is less clear, but qiskit and qpanda are consistently among the fastest packages (qsimcirq only supports single precision).
Figure 6: (a) Singlethread performance (double precision). (Top) Absolute wall clock time as a function of qubit number for packages that support double precision according to the package documentation. (Bottom) Performance ratio with respect to optimally performing package, here qiskit. Note that numbers above one correspond to a relative slow-down. (b) Scaling behavior of singlethread double precision performance by extracting $a$, $b$ as detailed in Fig. 4 for different quantum algorithms. (Top) Relative speed-down with respect to the reference package, qiskit. (Bottom) The red line indicates the expected scaling following $\log(2)$ as a function of the number of qubits. The paucity of data points in the large $N$-limit for the package myqlm_cpp limit the extraction of $a$ and $b$. $*{}^{*}$ In the an effort to minimze the errors on $a$ and $b$ for package braket, the final data point has been excluded while fitting.
@@
In Figs. 5(b) and  6(b) we quantify speed and computational overhead at large scales for the various packages using fit parameters $a$ and $b$. The package hybridq is excluded from this analysis due to a lack of data points in the large-$N$ limit, which results from the runtime limits we impose.
@@
Figure 9: In the case of QFT, $\braket{\sigma_{z}^{i}}=0$ at each site $i$, thereby allowing to sort the packages based on their precision performance, (a) single precision, (b) double precision. The red line marks the expected values for the precision set. In (a), smaller values point to a different precision used within the package compared to the precision supported in the documentation. For instance, packages like qiskit and pennylane document single precision settings but we note that the single precision settings turn out to be double precision.
@@
Three main types of limitations where found, which we classify as design limited, time limited and memory limited. Design limitation refers to the inability of the simulation package to go beyond a certain system size due to design decisions considered as a part of the development of the package. Design decisions range over a wide spectrum that include: the choice of the language used to develop the package, the choice of the core support libraries (for instance: numpy, scipy, AVX acceleration among others), design features to support multi-compute capabilities and so on. Time limited and memory limited are more subjective to the cluster on which the performance is being benchmarked. In our case, we limited the time to a day and the memory to 300 GB on the CPU node and 320 GB-GPU/900 GB-CPU on the GPU node.
@@
HiQ
$T(30),M(36)$
$T(30),M(36)$
HybridQ
$D(32)$
$D(32)$
$M(30),D(34)$
Intel-QS(cpp)
$T(28),M(36)$
$T(28),M(36)$
$T(32),M(36)$
$T(32),M(36)$
myQLM (py)
$T(28),D(34)$
myQLM (cpp)
$RA(22)$
$RA(22)$
$RA(22)$
$RA(22)$
Pennylane (py)
$T(28),D(34)$
$T(28),D(34)$
Pennylane (cpp)
$D(34)$
$T(32),D(34)$
$D(34)$
$T(32),D(34)$
$RE$
$M(32)$
Projectq
$T(32),M(34)$
$M(34)$
Qcgpu
$F(32),M(34)$
Qibo
$T(28),D(32)$
$T(28),D(32)$
$T(28),D(32)$
$T(28),D(32)$
Qibojit
$T(32),M(36)$
$T(30),M(36)$
$T(34),M(36)$
$T(34),M(36)$
$M(34)$
$M(32)$
Qiskit
$T(32)$
$T(32)$
$M(34)$
$T(34)$
$T(34)$
$T(32)$
QPanda
$T(32),M(34)$
$T(32),M(36)$
$NO$
Qrack
$T(30),M(36)$
$T(30),M(36)$
$T(34),M(36)$
$T(32),M(36)$
$M(32)$
$M(32)$
Qsimcirq
$D(34)$
$D(34)$
$M(34)$
Quest
$T(30),M(36)$
$T(30)$
$T(34),M(36)$
$T(34)$
$M(32)$
Qulacs
$T(32),M(36)$
$T(34),M(36)$
$M(30)$
Q++
$T(22)$
$T(22)$
SV-Sim
$D(30)$
$D(30)$
$D(30)$
Yao
$T(30),M(36)$
$T(30),M(36)$
$T(34),M(36)$
$T(32),M(36)$
$T(32),M(34)$
$M(32)$
@@
Data availability
The workflow used for generating the benchmarking results, including the singularity images, the quantum algorithms in the QASM format and sample job scripts can be found at the data repository [89]: https://zenodo.org/records/10376217. A web platform for online exploration and on-demand comparison across packages, tasks and hardware modalities can be found at https://qucos.qchub.ch.

### v2

@@ Introduction
The rapidly growing interest in quantum computers from both the academic and commercial side fuels the need for software packages that can simulate their operation and aid both hardware and algorithm development. These software environments generally model the quantum state evolution of an isolated set of qubits acted on by a set of well-defined quantum gate operations with an intrinsic error rate set only by the classical computer’s numerical precision. Some packages further support an open quantum systems simulation, which can also model external perturbations, thereby allowing for the simulation of actual, noisy quantum hardware at the expense of additional computational overhead. In all cases, the available computational hardware, the underlying mathematical formalism [1] and its particular software implementation place stringent limits on the system sizes that can be explored. With the exception of a few special cases, even High Performance Computing (HPC) platforms often cannot reach beyond a system size of about 50 simulated qubits before they run into resource limitations [2, 3, 4, 5]. Yet, even at moderate scales of a few tens of qubits, efficient simulation with reasonable execution times is a formidable challenge due to the vast amounts of data that have to be processed to faithfully store and manipulate a generic quantum state. With an amount scaling exponentially as $2^{\textrm{\#qubits}}$, memory and the transfer of data between processing and storage are at the core of the slow-down in computation speed and this is where HPC architecture and a tailored software implementation can make a difference.
@@
qiskit[7], cirq[8], qsimcirq[9], pennylane[10], pennylane lightning[10], qibo[11], qibojit[12], Intel QS[13], projectq[14], qrack[15], qpanda[16], qcgpu[17], quest[18], qulacs[19], qpp[20], SV-Sim[21], Yao[22], HiQ[23], HybridQ[24], Braket, myQLM, QuTiP[25], PyQuil[26], pytket[27], Microsoft QDK Simulator, Blueqat (qaqarot), Quantastica Toaster, HyQuas[28], MPIQulacs[29], JUQCS[30], Quimb[31], NVidia cuQuantum[32], Spinoza[33], QuantumFlow, QPlayer[34], Torchquantum[34], pyqtorch, QCompute,QX Simulator, Basiq[35], MIMIQ (QPerfect), Qristal, QCLAB++[36],quantum-gates [37], Qrisp
@@
qiskit, cirq, pennylane, qibo(jit), braket, hybridq, intel-qs, myqlm(py), qpanda, qsimcirq, quest, qulacs, q++, sv-sim, yao, quantum-gates, QuantumSim [38], NWQSim [21, 39], QuaC, QuTiP (PIQS), OpenQuantumTools.jl [40]
@@
qiskit, cirq, qrack, Stim [52], QuantumClifford.jl, PyZX [53], MIMIQ (QPerfect), pennylane
@@
First, there are statevector-based simulators that simulate the evolution of pure quantum states. Here, a statevector of $N$ qubits is represented by a $2^{N}$-sized vector of complex valued entries which are modified under the action of a formally $2^{N}\times 2^{N}$-sized unitary matrix operator. A more complex (and resource intensive) simulation is offered by the density matrix formalism, which further allows to simulate mixed states, which usually arise under the influence of noise and in the case of open quantum systems. Tensor networks offer an alternate representation of the quantum state. In the example of Matrix Product States (MPS) or Projected Entangled Pair States (PEPS) [68, 69] the wavefunctions are expressed in a compressed way using a set of local tensors to be contracted. The unitaries are represented as Matrix Product Operators (MPO) or Projected Entangled Pair Operators (PEPO), which are smaller in size than the above unitaries. However, with increasing entanglement in the system, the size of the tensors, referred to as the so-called bond dimension, is forced to grow to maintain an accurate approximation. Clifford algebra based simulators have been developed [70] that allow the access of qubit numbers on the order of a million and more. They are mainly used to investigate quantum error correction codes [52] and more recently random unitary circuits. Yet, these have been shown to not be universal for quantum computation [70] and adding non-Clifford operations to these simulations, while extending their range, rapidly makes them far less efficient. Both approaches are the leading methods in exploring large scale simulations at the edge of exact verifiability [71, 72].
Further, there are simulators that are tailored to a specific hardware architecture or application domain, aiming to emulate the corresponding platform exactly or being optimized for a specific use-case.
Table 2: List of benchmarked quantum simulation packages and supported features
Package Language Version ST MT GPU m-GPU MPI Shot Noise $\textbf{VQA}^{\dagger}$
SP DP SP DP SP DP SP DP
$\text{Braket}^{[k]}$ Python 1.38.1 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Cirq[8] Python 1.1.0 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
$\text{cuQuantum }^{[j]}$[73] Python 22.11.0 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[m]}$ $\checkmark^{[m]}$ $\checkmark^{[m]}$
HiQ[23] Python 0.0.1 $\checkmark$ $\checkmark$ $\checkmark^{[b]}$ $\checkmark^{[n]}$
HybridQ[24] Python 0.8.2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Intel-QS(cpp)[13] C++ 2.1.0 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[d]}$ $\checkmark^{[o]}$checkmark $\checkmark$
myQLM (py) Python 1.7.3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Pennylane(py)[10] Python 0.28.0 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[p]}$
Pennylane(cpp)[10] C++ 0.28.2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[c]}$ $\checkmark^{[b]}$
Projectq[14] Python 0.8.0 $\checkmark$ $\checkmark$ $\checkmark$
Qcgpu[17] Python 0.1.1 $\checkmark$ $\checkmark$
Qibo[11] Python 0.1.11 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Qibojit[12] Python 0.0.7 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Qiskit[7] Python 0.41.0 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[a]}$checkmark $\checkmark$ $\checkmark$
QPanda[16] Python 3.7.16 $\checkmark$ $\checkmark$ $\checkmark^{[h]}$checkmark $\checkmark$ $\checkmark$
Qrack[15] C++ 8.2.2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[q]}$ $\checkmark^{[r]}$
Qsimcirq[9] Python 0.15.0 $\checkmark$ $\checkmark$ $\checkmark^{[b]}$checkmark $\checkmark$ $\checkmark$
Quest[18] C 3.5.0 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[d]}$checkmark $\checkmark$
Qulacs[19] Python 0.5.7 $\checkmark$ $\checkmark$ $\checkmark^{[e]}$ $\checkmark^{[f]}$checkmark $\checkmark$ $\checkmark$
Q++[20] C++ 4.0.1 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SV-Sim[21] Python 1.0 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[g]}$ $\checkmark^{[d]}$checkmark $\checkmark^{[s]}$ $\checkmark^{[s]}$
Yao[22] Julia 0.8.6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark^{[d]}$ $\checkmark^{[d]}$checkmark $\checkmark$ $\checkmark$
ST - Singlethread, MT - Multithread, Shot - Shot simulation/Measurement simulation, Noise - Support for noisy simulations,
VQA - Variational Quantum Algorithms, SP/DP - Single/Double Precision, (m)-GPU - (multi) Graphical Processing Unit, MPI - Message Passing Interface,
We checkmark packages that offer modules with the statevector-optimizer integration,
@@
$[m]$ Available in CUDA Quantum that uses CuQuantum simulators, $[n]$ Derived from ProjectQ,
$[o]$ Measurements can be derived from probability of a qubit in a particular state, $[p]$ Fundamental design allows for the differentiation of
variational parameters, $[q]$ Support for shots in different providers, $[r]$ Not the typical density matrix support, $[s]$ Q#.
In this manuscript, we restrict our analysis to statevector simulators. Furthermore, our selection is based on additional criteria, such as, whether a package offers to exploit the computational power provided by HPC platforms, is actively maintained and developed (gauged in terms of release cycles) and provides access to relevant documentation (feature snippets, examples and possibly use-case tutorials). Table 2 presents our down-selected list of packages benchmarked in this work and their supported HPC capabilities [^1] While we list the MPI capability, we do not present MPI benchmarks in this work., in addition to the support for additional features like shot simulation, noise modelling and variational quantum algorithms.
@@
Figure 1: Workflow of the toolchain which takes four inputs: (i) Quantum Task, (ii) Simulation package, (iii) Compute capability of the HPC, (iv) Float precision. The task is input as the QASM instruction set which is processed by the translator to further generate the package specific instruction set. Post translation, other auxiliary file generation is triggered that creates the necessary job scripts, linker files that scope the translated run files into the environment of the container. In addition, a script to trigger the job submission onto the cluster is also generated. Finally, the performance characteristics either corresponding to the time or memory are collected for further analysis. () The translated files generated from OpenQASM are executed using the package installed on the container, while the pre- and post-processing steps are executed on the HPC node independent of the container.
@@
One of the first experiments that established claims of quantum advantage was the Random Quantum Circuits (RQC) sampling experiment implemented on the Sycamore chip [84]. In our benchmarks, we have used QASM files (specifically, the set marked as the EFGH pattern) supplied by the authors of the original experiment.
@@
QFT maps the state |$\ket{\psi}=\sum\limits_{i=0}^{M-1}\alpha_{i}\ket{i}$ where $M=2^{N}$ bitstrings (basis states) to |$\ket{\phi}=\sum\limits_{i=0}^{M-1}\beta_{i}\ket{i}$ where $\beta_{k}=\frac{1}{\sqrt{M}}\sum\limits_{m=0}^{M-1}\alpha_{m}\omega_{M}^{mk}$ with $k=0,1,2,3,...,M-1$ and $\omega_{M}=e^{\frac{2\pi i}{M}}$. The corresponding unitary is given by
$QFT_{N}=\begin{bmatrix}$\footnotesize{1}$&$\footnotesize{1}$&$\footnotesize{1}$&$\footnotesize{1}$&$\footnotesize{\ldots}$&$\footnotesize{1}$\\ $\footnotesize{1}$&$\footnotesize{$\omega_{M}$}$&$\footnotesize{$\omega_{M}^{2}$}$&$\footnotesize{$\omega_{M}^{3}$}$&$\footnotesize{\ldots}$&$\footnotesize{$\omega_{M}^{M-1}$}$\\ $\footnotesize{1}$&$\footnotesize{$\omega_{M}^{2}$}$&$\footnotesize{$\omega_{M}^{4}$}$&$\footnotesize{$\omega_{M}^{6}$}$&$\footnotesize{\ldots}$&$\footnotesize{$\omega_{M}^{2(M-1)}$}$\\ $\footnotesize{1}$&$\footnotesize{$\omega_{M}^{3}$}$&$\footnotesize{$\omega_{M}^{6}$}$&$\footnotesize{$\omega_{M}^{9}$}$&$\footnotesize{\ldots}$&$\footnotesize{$\omega_{M}^{3(M-1)}$}$\\ $\footnotesize{$\vdots$}$&$\footnotesize{$\vdots$}$&$\footnotesize{$\vdots$}$&$\footnotesize{$\vdots$}$&$\footnotesize{$\ddots$}$&$\footnotesize{$\vdots$}$\\ $\footnotesize{1}$&$\footnotesize{$\omega_{M}^{M-1}$}$&$\footnotesize{$\omega_{M}^{2(M-1)}$}$&$\footnotesize{$\omega_{M}^{3(M-1)}$}$&$\footnotesize{\ldots}$&$\footnotesize{$\omega_{M}^{(M-1)(M-1)}$}$\\ \end{bmatrix}.$
@@
$QFT(\ket{\varphi})$ $=\frac{1}{\sqrt{M}}\Big{\{}\bigg{(}\ket{0}+e^{2\pi i[0.\varphi_{n-1}]}\ket{1}\bigg{)}$
@@
where $[0.\varphi_{0}\varphi_{1}...\varphi_{l}]=\sum\limits_{p=0}^{l}\varphi_{p}2^{-p}$. The $\sigma_{z}$ expectation value of Eq. (9) at any site $i$ will be equal to zero, which we use to validate the computations and order the benchmarked packages based on their precision via their proximity to zero, as detailed in a later section.
@@ Results
In this part of our study, we configure the toolchain to limit the available compute capability to a single computational thread. The aim is to investigate the effect of numerical precision (single vs. double) on both the wall-clock time scaling as well as the maximum system size that can be simulated. While this difference of 32 bit vs. 64 bit representation will not allow for a much larger number of qubits (a factor two reduction in total memory), less data will have to be processed at every step, allowing for speed-ups depending on the package internals.
@@
We start by noting that in the top row of Figs. 5(a) and 6(a), we observe that some packages exhibit almost straight lines for the entire range of system sizes, while other packages have a rather small slope at lower numbers of qubits and only cross over to the steeper slope for qubit numbers of about 20. This leads to the paradoxical situation that the fastest package in Fig. 6(a) at the largest qubit numbers (qiskit) is among the slowest packages for eight qubits  [^2] A similar observation holds for hybridq in the single precision case, although it is never the fastest package, even though it gets quite close..
Among the packages that support single precision, shown in Fig. 5(a), we can clearly identify that the package qsimcirq performs best in the large-$N$ limit. For double precision resolution, shown in Fig. 6(a), the best performing package is less clear, but qiskit and qpanda are consistently among the fastest packages (qsimcirq only supports single precision).
Figure 6: (a) Singlethread performance (double precision). (Top) Absolute wall clock time as a function of qubit number for packages that support double precision according to the package documentation. (Bottom) Performance ratio with respect to optimally performing package, here qiskit. Note that numbers above one correspond to a relative slow-down. (b) Scaling behavior of singlethread double precision performance by extracting $a$, $b$ as detailed in Fig. 4 for different quantum algorithms. (Top) Relative speed-down with respect to the reference package, qiskit. (Bottom) The red line indicates the expected scaling following $\log(2)$ as a function of the number of qubits. The paucity of data points in the large $N$-limit for the package myqlm_cpp limit the extraction of $a$ and $b$. In the an effort to minimze the errors on $a$ and $b$ for package braket, the final data point has been excluded while fitting.
@@
In Figs. 5(b) and  6(b) we quantify speed and computational overhead at large scales for the various packages using fit parameters $a$ and $b$. As we are unable to obtain reliable fit parameters for the package hybridq due to the paucity of data in the large-$N$ limit, we exclude it from the comparison.
@@
Figure 9: Accumulated errors in QFT for N=16 with regards to baseline expectation in (a) single precision and (b) double precision settings. Summed up deviations from the expected $\braket{\sigma_{z}^{i}}=0$ are shown for each setting (red line) and packages sorted correspondingly. We note that in (a) packages like qiskit and pennylane document single precision settings but are found to be running double pression, instead.
@@
Three main types of limitations were found, which we classify as design limited, time limited and memory limited. Design limitation refers to the inability of the simulation package to go beyond a certain system size due to design decisions considered as a part of the development of the package. Design decisions range over a wide spectrum that include: the choice of the language used to develop the package, the choice of the core support libraries (for instance: numpy, scipy, AVX acceleration among others), design features to support multi-compute capabilities and so on. Time limited and memory limited are more subjective to the cluster on which the performance is being benchmarked. In our case, we limited the time to a day and the memory to 300 GB on the CPU node and 320 GB-GPU/900 GB-CPU on the GPU node.
@@
HiQ $T(30),M(36)$ $T(30),M(36)$
HybridQ $D(32)$ $D(32)$ $M(30),D(34)$
Intel-QS(cpp) $T(28),M(36)$ $T(28),M(36)$ $T(32),M(36)$ $T(32),M(36)$
myQLM (py) $T(28),D(34)$
myQLM (cpp) $RA(22)$ $RA(22)$ $RA(22)$ $RA(22)$
Pennylane (py) $T(28),D(34)$ $T(28),D(34)$
Pennylane (cpp) $D(34)$ $T(32),D(34)$ $D(34)$ $T(32),D(34)$ $RE$ $M(32)$
Projectq $T(32),M(34)$ $M(34)$
Qcgpu $F(32),M(34)$
Qibo $T(28),D(32)$ $T(28),D(32)$ $T(28),D(32)$ $T(28),D(32)$
Qibojit $T(32),M(36)$ $T(30),M(36)$ $T(34),M(36)$ $T(34),M(36)$ $M(34)$ $M(32)$
Qiskit $T(32)$ $T(32)$ $M(34)$ $T(34)$ $T(34)$ $T(32)$
QPanda $T(32),M(34)$ $T(32),M(36)$ $NO$
Qrack $T(30),M(36)$ $T(30),M(36)$ $T(34),M(36)$ $T(32),M(36)$ $M(32)$ $M(32)$
Qsimcirq $D(34)$ $D(34)$ $M(34)$
Quest $T(30),M(36)$ $T(30)$ $T(34),M(36)$ $T(34)$ $M(32)$
Qulacs $T(32),M(36)$ $T(34),M(36)$ $M(30)$
Q++ $T(22)$ $T(22)$
SV-Sim $D(30)$ $D(30)$ $D(30)$
Yao $T(30),M(36)$ $T(30),M(36)$ $T(34),M(36)$ $T(32),M(36)$ $T(32),M(34)$ $M(32)$
@@
Data and Code availability
The entire source code for the QASM translator, toolchain, plotting/data analysis and auxillary files as well as the extracted data can be found at https://huggingface.co/spaces/amitjamadagni/qs-benchmarks/tree/main. Additionally, the workflow used for generating the benchmarking results, including the singularity images, the quantum algorithms in the QASM format and sample job scripts can be found at the data repository [92]: https://zenodo.org/records/10376217. An interactive web platform for online data exploration and on-demand comparison across packages, tasks and hardware modalities can be found at https://qucos.qchub.ch.